Parth (parthkohli):

What is the remainder when \(2^{64}\) is divided by \(7\)?

Parth (parthkohli):

I get a pattern!

mathslover (mathslover):

2^2 (which equals 4) divided by 7 leaves remainder 4. 2^3 (which equals 8) divided by 7 leaves remainder 1. 2^4 (which equals 16) divided by 7 leaves remainder 2. 2^5 (which equals 32) divided by 7 leaves remainder 4. 2^6 (which equals 64) divided by 7 leaves remainder 1.

Parth (parthkohli):

2mod7 = 2 2^2 mod 7 = 4 2^3 mod 7 = 1 and it is repeating!

mathslover (mathslover):

i think it is : 1 ..

mathslover (mathslover):

64/7 =remainder = 1 ..

Parth (parthkohli):

For every 2 powers, it gets 2 as the remainder.

Parth (parthkohli):

Why are you dividing the exponent? :P

Parth (parthkohli):

For every four*

Parth (parthkohli):

And 64 is divisible by 4, so it's 2?

Parth (parthkohli):

Hmm, it isn't. :/

Parth (parthkohli):

2^8 mod 7 = 4 So that statement was completely false.

OpenStudy (anonymous):

Umm I don't know about a proof but you have a repeating pattern with a period of 3, predicting that the remainder will be the same as for 2^4, that is, 2 Python: >>> 2**64 18446744073709551616L >>> 2**64 % 7 2L Seems right.

mathslover (mathslover):

yes the answer is 4.

mathslover (mathslover):

oops sorry i meant the answer is 2

OpenStudy (anonymous):

Scared me.

mathslover (mathslover):

:D ..

Parth (parthkohli):

Hmm, but that thing I said didn't satisfy for \(2^8\).

mathslover (mathslover):

2^8/7 = remainder = 2

OpenStudy (anonymous):

>>> 2**8 % 7 4

mathslover (mathslover):

yes .. remainder = 4 .. but why m i getting 2 by that way?

mathslover (mathslover):

256/7 = remainder = 4

mathslover (mathslover):

The answers follow this pattern: 2 divided by 7 leaves remainder 2. 2^2 (which equals 4) divided by 7 leaves remainder 4. 2^3 (which equals 8) divided by 7 leaves remainder 1. 2^4 (which equals 16) divided by 7 leaves remainder 2. 2^5 (which equals 32) divided by 7 leaves remainder 4. 2^6 (which equals 64) divided by 7 leaves remainder 1. We should stop as soon as we notice that the cycle will repeat itself forever in this pattern: [2, 4, 1]. Every third remainder is the same. (From here on out, "remainder" always means "remainder after we divide by 7.") Since every third remainder is the same, we should look at the remainder when the power is a multiple of 3. The remainders of 2^3 and 2^6 are 1. Thus, the remainder of 2 raised to a power that is any multiple of 3 is 1.

Parth (parthkohli):